OpenStudy (walkingcontradiction):

I need help. WILL AWARD MEDAL AND FAN... maybe a fan for a fan :D Jk, well you could ;D

pooja195 (pooja195):

Hi Welcome to openstudy c: Please post your question ^-^

OpenStudy (md152727):

:)

pooja195 (pooja195):

Is this a test?

OpenStudy (walkingcontradiction):

It's a practice test.

OpenStudy (walkingcontradiction):

Doesn't count for a grade FLVS :/

pooja195 (pooja195):

I'm not sure about this one sorry :(

OpenStudy (walkingcontradiction):

Do you know anyone that can help cuz I'm stuck big time

pooja195 (pooja195):

Yeah i posted a link in math chat hopefully someone will help

OpenStudy (walkingcontradiction):

Yea, I saw thanks :) I feel so popular with all the views ;p

OpenStudy (md152727):

@Hero @ParthKohli @dan815 @Abhisar @AlexandervonHumboldt2 @Compassionate @jigglypuff314

jigglypuff314 (jigglypuff314):

not entirely sure... if it proportional, then... 16/24 = (BD+16)/(24+36) and 27/36 = (BF+27)/(24/36) I think... but I remember getting these sorts of problems wrong myself back in the day ^_^"

OpenStudy (walkingcontradiction):

So what is it? 24 As we can see I suck at geometry @jigglypuff314

jigglypuff314 (jigglypuff314):

well I tried converting geometry stuff into algebra for you ^_^ take a look at the two equations I gave in my previous comment, can you solve them?

OpenStudy (walkingcontradiction):

Nope I Suck at math :(

OpenStudy (walkingcontradiction):

@jigglypuff314

jigglypuff314 (jigglypuff314):

16/24 = (BD+16)/(24+36) 2/3 = (BD + 16) / (60) and 27/36 = (BF+27)/(24+36) 3/4 = (BF + 27) / (60) \[\frac{ 2 }{ 3 } = \frac{ BD + 16 }{ 60 }\]\[\frac{ 3 }{ 4 } = \frac{ BF + 27 }{ 60 }\]

OpenStudy (walkingcontradiction):

So the answer is... :D

jigglypuff314 (jigglypuff314):

try to solve for BD and BF

OpenStudy (walkingcontradiction):

OMFG I CANT... IM LITERALLY CRYING

OpenStudy (walkingcontradiction):

@jigglypuff314

OpenStudy (walkingcontradiction):

@jigglypuff314